OpenStudy (anonymous):

7-4x(486)/what-1=724

OpenStudy (anonymous):

um aigth so i kind of get what the question is asking but can you be a little more clear on the question

OpenStudy (anonymous):

7-4x*486/x-1 =724…is this what you are asking

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aight hold on i'm havin sumone checkmy answer real quick

OpenStudy (anonymous):

Aight I got one more question does the equation look like this: \[7-4x\frac{ 486 }{ x-1 } = 724\] or \[\frac{ 7-4x(486) }{ x-1 } =724\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wich one?

OpenStudy (anonymous):

2nd

OpenStudy (anonymous):

well i got 731over 2668 but i'm not 100 on it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

aight dnt foget to give me one of those medal thingys lol